As the pressure of a gas at 101.3 kpa is changed to 50.65 kpa at constant temperature the volume of gas
1
Expert's answer
2015-11-23T14:31:50-0500
Under Boyle and Marriott's law, at a constant temperature work of pressure of gas upon its volume is a constant
P1*V1 = P2*V2
Then V2 = P1V1/P2 = V1 * 101.3/50.35 = 2V1

Answer: Volume will increase twice